### v3

# Pretraining Frame Preservation in Autoregressive Video Memory Compression

Abstract
We present a neural network structure to compress long videos into short contexts, with an explicit pretraining objective to preserve the high-frequency details of single frames at arbitrary temporal positions. The baseline model can compress a 20-second video into a context at about 5k length, where random frames can be retrieved at perceptually preserved appearances. Such pretrained models can be directly fine-tuned as memory encoders for autoregressive video models, enabling long history memory with low context cost and relatively low fidelity loss. We evaluate the framework with ablative settings and discuss the trade-offs of possible neural architecture designs.
Figure 1: This video was generated autoregressively second-by-second with full history context (without cutting out any history frames). The 20+ second history was compressed into a $\sim$5k context length and processed by an RTX 4070 12GB.
1 Introduction
Storytelling capability, narrative coherence, and context quality are increasingly important in the video generation community. The latest commercial models like Sora2 [46], Veo3.1 [20], and Kling2.5 [37] enable storyboard creation-based, scene planning, and dynamic camera shots (e.g., in the style of TikTok or YouTube Shorts). Recent academic models have also focused on long video streaming [72, 27], scene planning [22, 26, 83, 86, 87], and video context consistency [28, 31, 42]. Among these approaches, autoregressive models are a key paradigm for video storytelling, supporting native video continuation and storyboard streaming. Autoregressive models view video history as context, and they face unique challenges when handling long-form content like in-the-wild videos and movies.
The trade-off between context quality and length is a typical and critical problem in video context modeling. A naive sliding window, which cuts off all distant-enough frames, maintains a fixed context length but loses long-range history. Various compression approaches have been extensively studied; e.g., highly compressed VAEs like LTXV [23] and DC-AE [10] show that it is possible to compress pixel-space videos into more compact contexts, and hybrid approaches [1, 32, 88] like FramePack [78] show that compression can happen at multiple levels, albeit at the cost of high-frequency image details. Another strategy is to preserve the context length but reduce computation (e.g., using sparse [64, 77, 80, 65] or linear [5, 68, 58, 11, 74, 33] attention), though the linear layer cost remains an overhead for training and bidirectional inference. More specialized methods like token merging [4, 3] also demonstrate that a higher merging rate leads to greater detail loss. These research efforts implicitly converge on the observation that context quality and length are governed by a fundamental trade-off where the optimal balance has not yet been reached.
This observation leads to a wide range of open questions. Is there an explicit correlation between context length and quality, and how can we model it? Should context quality be measured by object, scene, or facial consistency, or potentially, by full-frame reconstruction quality? To achieve a perceptually acceptable quality from a human perspective, what is the maximum level of compression, and where is the sweet spot? How can such a balance be found? Do different tasks and objectives pose varied requirements for this trade-off? Do spatial and temporal compression have different levels of importance in various tasks? Is there a marginal effect of context length that makes a sufficiently long context actually redundant?
To study these questions, this paper targets an explicit objective: a learning task to directly compress video history frames into a minimized context length while maximizing frame retrieval quality. This setting allows us to explicitly study the correlation between context length and reconstruction quality using both quantitative metrics and human perceptual evaluations. We can thereby independently study the quality, influence, and cost of the compression model before moving on to the actual autoregressive video generation. This white-box feedback eases cost management and contributes to the explainability of separate encoding components in autoregressive video models.
To this end, one challenge is that learning to compress and transform a long video context causes computational bottlenecks. We observe that a strong indicator of a video compression mechanism’s ability to preserve contextual details is its capacity for high-quality frame reconstruction at arbitrary temporal locations. This leads to a randomized approach: we first pretrain a memory encoder to reconstruct frames at random time positions, and then finetune an autoregressive video model using the compression model as a history memory encoder. Pretraining with millions of videos, we show that this method reduces the video model training cost by a large margin and facilitates a stronger neural representation for detailed frame retrieval.
Experiments show that the pretrained memory encoder can, by itself, reconstruct history frames at high quality (measured by PSNR/SSIM [60]). When connected to video diffusion models for autoregressive finetuning, the encoder provides long-range consistency. This framework achieves a balanced and optimized point in the context length-quality trade-off, leading to a practical autoregressive video model with a relatively short context length and perceptually high consistency based on its long-range history.
In summary, we (1) discuss the trade-off between context length and quality in autoregressive video generation; (2) propose to pretrain an explicit compression task for frame retrieval for improved autoregressive video memory; (3) establish an autoregressive system to fine-tune the pretrained compression model as a history memory encoder for DiTs; (4) conduct extensive experiments, including quantitative, qualitative, and ablation studies, to analyze the influence of different compression designs; and (5) provide pretrained models for practical video generation and editing applications.
2 Related Work
Autoregressive video diffusion and long videos. Diffusion has become the driving force behind video synthesis, where a central challenge is length scaling. Training-free length‑extension methods [48, 43, 44, 84, 49] reschedule noise or re-balance temporal frequency to stretch pre-trained models beyond their training horizon. A complementary thread blends diffusion with causal prediction: Diffusion Forcing [7] and HistoryGuidance [52] enable variable-horizon conditioning and stable long rollouts by noise injection. These approaches are adapted in industrial systems such as SkyReels‑V2 [8] and Magi-1 [55]. Additionally, StreamingT2V [24] augments existing models with short- and long-term memory with randomized blending to generate longer videos. FAR [21] studies long‑context AR with flexible RoPE [53] decay and mixed short/long windows, while CausVid [72] distills a bidirectional teacher to a few‑step causal generator. StreamDiT [35] combines multi-step distillation with a moving frame buffer and mixed partition training to generate results in real-time. To mitigate error accumulation during AR generation, Self‑Forcing [27] simulates AR rollout during training, while its extensions [12, 41, 69] further improve length generalization.
Context learning and compression. Long-context persistence is crucial as we extend video generation beyond a few seconds. One option is retrieval, which grounds generation in a persistent state. WorldMem [67] and Context-as-Memory [73] augment AR video world models with FoV-based history retrieval, while VMem [39] indexes past views by surfels to retrieve the most relevant viewpoints. Memory Forcing [25] pairs geometry-indexed spatial memory with tailored training regimes to balance exploration and revisits. Beyond fixed retrieval rules, Mixture-of-Contexts [6] learns a dynamic sparse attention context router so that tokens attend only to the most salient chunks. Similarly, MoGA [30] and Holocine [45] also propose token-level sparse attending policies. Pack-and-Force [63] instead proposes a learnable context semantic retriever. Orthogonal to retrieval, compression of history turns an unbounded context into a compact state. FramePack [78] compresses prior frames into a fixed‑size latent “packed” context. Captain Cinema [66] uses a similar compression for keyframes. StateSpaceDiffuser [50] and Po et al. [47] swap quadratic attention for recurrent states to maintain long‑term memory. TTTVideo [13] and LaCT [81] use light MLP test-time training layers as a learned context representation.
Efficient video diffusion designs. As we scale video generation to long horizons, large context windows are bottlenecked, driving a wave of efficient computational designs. Kernel advances such as FlashAttention [14, 15] improve throughput. Static or hardware-friendly sparse patterns include sliding/tiling 3D windows [80], radial spatiotemporal masks [40], and training-free head/pruning heuristics [64, 70]. Dynamic or learned pruning/routing further select salient token pairs or blocks [62], coarse-to-fine sparse token selection [79], Sage/SpargeAttention families [76, 75, 77], blockified routing with cached search [65], and progressive block carving [82]. Another line of work leverages compressing the latent space or token sequence: token merging and patch scaling [2, 38], compact/variable-rate tokenizers [1], highly compressed latent space [23], or multiscale pyramids with re-noising [32]. SANA [68] introduced a linear‑attention Diffusion Transformer for images, while SANA‑Video [9] extends this with block‑linear attention and a constant‑memory KV cache.
Besides, this approach presents distinctions with VAEs: VAEs have bottlenecks with 4, 16, or 64 channels, whereas this method directly works at DiT inner channels like 3072 or 5120 with a different feature scope; VAEs have training overheads to process full contexts, while this method works with lightweight partial overheads (retrieval objective).
3 Method
We pretrain a video compression model with the objective to retrieve frames at arbitrary time positions, and simultaneously preserve frame details (Section 3.1). The pretrained compression model is then finetuned as a memory encoder for autoregressive video diffusion models (Section 3.2).
Preliminaries. We clarify the notation of a typical autoregressive video diffusion model. Unless otherwise noted, all “frames”, “pixels”, etc., refer to latent concepts. We consider the typical Diffusion Transformers (DiTs) like Wan [56] and HunyuanVideo [36] with flow match scheduling. Rectified-flow models map noisy latents $\bm{X}_{t_{i}}\in\mathbb{R}^{T\times H\times W\times C}$ with
$\bm{X}_{t_{i}}=(1-t_{i})\bm{X}_{0}+t_{i}\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}\big(\mathbf{0},\mathbf{I}\big),$ (1)
from clean latents $\bm{X}_{0}$ where $t_{i}\in(0,1]$ is the diffusion timestep. The autoregressive model typically takes the history $\bm{H}\in\mathbb{R}^{T_{h}\times H\times W\times C}$ as conditioning. To learn the autoregressive generator $\bm{G}_{\theta}(\cdot)$, the learning objective is
$\vskip-3.0pt\mathbb{E}_{\bm{X}_{0},\bm{H},\bm{c},\bm{\epsilon},t_{i}\sim\mathcal{L}(0,1)}\bigg\|(\bm{\epsilon}-\bm{X}_{0})-\bm{G}_{\theta}\big(\bm{X}_{t_{i}},t_{i},\bm{c},\bm{H}\big)\bigg\|_{2}^{2},$ (2)
where $\bm{c}$ is a set of conditions such as text prompts (a single prompt or streamed from a storyboard), and $t_{i}\sim\mathcal{L}(0,1)$ is the shifted logit-normal distribution [18] for flow matching.
3.1 Pretraining Memory Compression Model
Figure 2: Pretraining of memory compression models. The memory compression model has to compress long videos (e.g., 20 seconds) into short contexts (e.g., of length 5k). The objective of the pretraining is to retrieve frames with high-frequency details in arbitrary history time positions.
Consider a typical case where we have a history of a 60-second video and want to generate the next 1 second, in a typical setting of 480p, 24fps using standard Hunyuan/Wan VAE and patchifying. The history context length becomes $832/16\times 480/16\times 60\times 24/4=561,600$, which creates an inference bottleneck for consumer-level devices that have limited GPU memory (and a training bottleneck for clusters even with causal computation or KV-cache designs). A naive sliding window will fail to model long-range consistency. Typical overhead reduction like token merging, downsampling, larger patch sizes, compact VAEs, etc., come with costs of losing image detail at various levels.
We observe that a well-suited indicator of a video compression mechanism’s ability to preserve contextual details is its capacity for high-quality frame retrieval at arbitrary time positions. For high compression rates, perfect retrieval becomes impractical, leading to a goal to maximize the retrieval quality of arbitrary frames.
We formulate this goal in a technical form. Considering that a compression process $\phi(\cdot)$ receives a history $\bm{H}$ of $T_{h}$ frames and compresses it into a context $\phi(\bm{H})$, given an arbitrary set of frame indices $\bm{\Omega}\subseteq\mathbb{N}$, and denoting a indexed retrieval process as $\phi^{-1}(\cdot,\bm{\Omega})$, we want to minimize the feature distance $\|\bm{H}_{\bm{\Omega}}-\phi^{-1}(\phi(\bm{H}),\bm{\Omega})\|_{2}^{2}$ for optimized reconstruction.
As shown in Fig. 2, the frame selection $\bm{\Omega}$ and retrieval $\phi^{-1}(\cdot,\bm{\Omega})$ can be established as an autoregressive video diffusion framework. We sample random videos as the history, and then pick random frames $\bm{\Omega}$: in particular, we randomly keep a subset of frames at indices $\bm{\Omega}$ unchanged and then mask all remaining frames. Various types of masks can be utilized like range scaling, shifting, etc. Herein, we use a noise-as-mask method: add latent noise levels from $\mathcal{L}(0.2,1)$. After that, we clone the picked clean frames as the diffusion target, so that the diffusion system will try to reconstruct the target frames at arbitrary positions. This process can also be written in line with Eq. (2) as
$\vskip-3.0pt\mathbb{E}_{\bm{H},\bm{\Omega},\bm{c},\bm{\epsilon},t_{i}}\bigg\|(\bm{\epsilon}-\bm{H}_{\bm{\Omega}})-\bm{G}_{\theta}\big((\bm{H}_{\bm{\Omega}})_{t_{i}},t_{i},\bm{c},\phi(\bm{H})\big)\bigg\|_{2}^{2},$ (3)
where $\bm{G}_{\theta}(\cdot)$ can be a trainable video diffusion model like WAN or HunyuanVideo with LoRAs.
When the output length of the compression model $\phi(\cdot)$ is short, e.g., about only 5k length for every 20 seconds, the retrieval task becomes highly difficult, making it well-suited for pretraining strong context compression models using millions of in-the-wild videos. We also point out the importance of the randomness in frame indices $\bm{\Omega}$ selection: if the frames are not randomly selected, e.g., only training with fixed starting or ending positions, the framework will learn a cheating solution, e.g., using all 5k context to encode a few ending frames while ignoring all other frames. Besides, note that directly training to retrieve all frames is impractical with minute-level data, even for highly efficient infrastructures.
Figure 3: Architecture of memory compression model. We use 3D convolution, SiLU, and attention to establish a lightweight neural structure as the baseline compression model. Different alternative architectures (e.g., various channels, full transformer, etc.) are possible and will be discussed in ablation.
Network architecture. As shown in Fig. 3, we present a lightweight baseline architecture for the compression model $\phi(\cdot)$. Instead of building the compressed representation from scratch, we reuse the DiT’s context representation: we first downsample a high-resolution high-fps video into a low-resolution low-fps one and process it with the VAEs and DiT’s patchifiers and first-layer projections. After that, we encode the high-resolution original video into a residual-enhancing vector and add it to the context vector. Note that the feature addition happens after the DiT’s first-layer projection, which means the encoder does not go through the VAE’s 16 channel bottleneck but directly outputs at inner channels (like 3072 for WAN-5B) for best fidelity. The baseline encoder uses 3D convolutions as starting layers and ends with attention layers. Besides, it is also possible to connect the DiT with cross-attentions from the last hidden states of the encoder for another refinement pass, but the quality gain is marginal and will be covered in ablations.
3.2 Finetuning Video Diffusion Model
Figure 4: Finetuning autoregressive video models. We illustrate the finetuning and inference of the final autoregressive video models. The pretraining of the memory compression model is finished before the finetuning.
With the pretrained memory compression model $\phi(\cdot)$, we can form autoregressive video generation systems by finetuning video diffusion models like WAN (with LoRAs) and the pretrained compression model as history memory encoders. This will lead to an autoregressive video model with long history windows (e.g., >> 20s), short history context length (e.g., about 5k), and explicitly optimized frame retrieval quality. The diffusion can be written in line with Eq (2) as
$\vskip-3.0pt\mathbb{E}_{\bm{X}_{0},\bm{H},\bm{c},\bm{\epsilon},t_{i}}\bigg\|(\bm{\epsilon}-\bm{X}_{0})-\bm{G}_{\theta}\big(\bm{X}_{t_{i}},t_{i},\bm{c},\phi(\bm{H})\big)\bigg\|_{2}^{2},$ (4)
where the process is as shown in Fig. 4-(a). The inference follows autoregressive generation by repeatedly concatenating the generated sections into the history as shown in Fig. 4-(b). Note that since the baseline encoder (Fig. 3) is almost fully convolutional, the compressed history can also be concatenated on-the-fly without recomputation.
Figure 5: Visual comparison on compression reconstruction. We present reconstruction results after the pretraining using different possible neural structures and various compression settings. * The “Large Patchifier” is technically equivalent to [78].
4 Experiment
4.1 Experimental Details
[…]
Hyper-parameters. When constructing the encoders of the memory compression models, we use the denotation $H\times W\times T$ to mark the compression rates, e.g., $4\times 4\times 2$ means compressing the latent width 4 times, height 4 times, and temporal length 2 times. Note that this compression rate is on top of the latent space and can be multiplied by the latent rate and patchifier rate to compute the final compression rate against the pixel space. When building the 3D convolution layers, we follow the rule to always first reduce the temporal dimension and then the spatial dimension (using strides in convolutions). The hidden channels of the encoder follow the pattern $64\rightarrow 128\rightarrow 256\rightarrow 512$ and are then always kept at $512$ (e.g., in the final attention layer) before the last $1\times 1$ convolution, which projects the feature to the same dimensions (like 3072, 5120, etc.) as the DiT context.
Base models. All experiments are conducted on base models of HunyuanVideo [36] and Wan [56] family. For Wan 2.2, the involved models are the 5B model and the 14B high noise model. Only the high noise model is used to simplify experiments (similar to the design choice of WanAnimate).
Data preparation. The dataset is prepared with about 5 million internet videos from a range of websites. About half of the data is vertical Short-style videos and the rest are common horizontal videos. The data is cleaned for quality, and then the high-quality part is captioned with Gemini-2.5-flash VLM and the remainder is processed with local VLMs like QwenVL [57]. The generated captions are in storyboard format with timestamps. When training the autoregressive models, only the prompt with the nearest timestamp is received by the model.
Storyboards. If not noticed, all storyboard prompts in this paper are written by Gemini-2.5-pro. We also tested open source alternatives like Qwen [57] and the quality difference is negligible. See also the supplement for instructions.
Test samples. The test set includes 1000 storyboard prompts written by Gemini-2.5-pro and 4096 unseen videos independent from the training dataset.
4.2 Ablation Study
Table 1: Quantitative results on compression structures. We show numerical tests with different ablative compression architectures. * The “Large Patchifier” is technically equivalent to [78].
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Large Pachifier* ($\times4$$\times$2) 12.93 0.412 0.365
Only LR ($\times4$$\times$2) 15.21 0.472 0.212
Without LR ($\times4$$\times$2) 15.73 0.423 0.198
Proposed ($\times4$$\times$2) 17.41 0.596 0.171
Proposed ($\times2$$\times$2) 19.12 0.683 0.152
Proposed ($\times2$$\times$4) 18.63 0.637 0.153
Proposed ($\times2$$\times$1) 20.19 0.705 0.121
Compression model architecture. As shown in Fig. 5, we compare the image reconstruction details of different architectures under various compression rates, focusing on the performance of Only LR, Without LR, Large Patchifier, and our proposed method. The Only LR variant keeps only the low-resolution branch in Fig. 3, while Without LR removes it, leaving only the high-resolution branch. The Large Patchifier method increases the compression rate by enlarging the DiT’s patchifying projection kernel without introducing new layers, equivalent to FramePack [78] and we directly use its implementation. The results indicate that the Only LR and Without LR variants produce reconstructions that differ significantly from the original images. The Large Patchifier method yields substantial structural changes. The proposed method, even at a high compression rate such as $4\times 4\times 2$, can still preserve much of the image appearance. A lower compression rate, such as $2\times 2\times 1$, provides the best detail preservation, albeit at the cost of a longer context length.
The quantitative results are presented in Table 1. These results show that this method achieves relatively higher performance on metrics like PSNR and SSIM. Furthermore, even at a higher compression rate of $4\times 4\times 2$, this method effectively preserves the original image structure.
Figure 6: Influence of memory compression model pretraining. We present results with or without the pretraining of the memory compression model. The input is the same 20-second history video, and we visualize the middle frames in the output shots.
Influence of pretraining. We demonstrate in Fig. 6 the impact of pre-training by comparing models trained with and without it. We keep the model architecture and the 20-second history input identical. One model is trained from randomly initialized encoder weights, while the other is fine-tuned using the pretrained memory encoder. Both models are trained with 100k sufficient steps. The results show a marked difference. With pre-training, the model maintains strong temporal consistency, leading to notable improvements in facial features, clothing, global video style, storytelling plots, and the coordination of camera movements. In contrast, the model trained without pre-training sometimes fails to attend to relevant frames in the history, resulting in inconsistencies.
Error accumulation (drifting). The problem of error accumulation highly depends on the training dataset. Surprisingly, we find that when trained on videos with dense camera shot shifts, like Short-style videos, the error accumulation problem seems to be absent or invisible. For other types of videos requiring long continuation over single shots, multiple design choices [78, 27, 7] are discussed in the supplement.
4.3 Qualitative Results
Figure 7: Qualitative results on storyboards. We present results by streaming prompts from storyboards. A storyboard is a set of prompts where each prompt covers a certain number of frames. The storyboards can be written by external language models.
Storyboard results. As shown in Fig. 7, we demonstrate that our model can handle a diverse range of prompts and storyboards while maintaining consistency across characters, scenes, objects, and plotlines.
4.4 Quantitative Results
Metrics. We include several metrics for video evaluations from VBench [29], VBench2 [85], etc., with some modifications. Cloth: The cloth consistency by asking VLM questions. We upgrade VBench2 [85]’s LLaVA to Gemini-2.5-pro; Instance: The object-level quality by asking VLM questions about unnatural or inconsistent instances. We upgrade VBench2 [85]’s Qwen to Gemini-2.5-pro; Identity: The facial similarity using ArcFace [16] with face detection RetinaFace [17]; Clarity: The MUSIQ [34] image quality predictor trained on SPAQ [19]. This metric evaluates artifacts like noises; Aesthetic: The LAION aesthetic predictor [51], measuring the aesthetic values perceived by a CLIP estimator; Dynamic: The RAFT [54] modified by VBench to estimate the degree of dynamics. Note that multi-shot videos intuitively have a special RAFT flow between shots, but in practice, RAFT seems to output near-zero values between shots, making the difference negligible from regular videos; Semantic: The video-text score computed by ViCLIP [59]. This metric measures the overall semantic consistency and prompt alignments.
Table 2: Quantitative results on video content consistency. We present the performance of different possible architectures on content consistency. The “1p” in Qwen means using 1 image as an image model input. Methods with severe artifacts are excluded from human ELO scores.
Human Object User Study
Method Cloth $\uparrow$ Identity $\uparrow$ Instance $\uparrow$ ELO $\uparrow$
WanI2V [56] + QwenEdit [61] (1p) 94.10 68.45 85.21 /
WanI2V [56] + QwenEdit [61] (2p) 95.09 68.22 91.19 1198
WanI2V [56] + QwenEdit [61] (3p) 94.28 67.30 80.34 /
Only LR ($\times4$$\times$2) 91.98 69.22 85.32 1194
Without LR ($\times4$$\times$2) 89.64 67.41 82.85 /
Without Pretrain ($\times4$$\times$2) 87.12 66.99 81.13 /
Proposed ($\times4$$\times$2) 96.12 70.73 89.89 1216
Proposed ($\times2$$\times$2) 96.71 72.12 90.27 1218
Consistency analysis. We compare various architectures of our method, including different compression rates and ablations with or without pre-training and the LR branch. We also compare against a common baseline that combines Wan-2.2-I2V [56] with Qwen-Image-Edit-2509 [61]. This baseline first generates a series of initial frames with an image model, animates them into video clips, and then concatenates the clips. We test this baseline with 1, 2, and 3 images for Qwen-Edit as history contexts (since the model supports at most 3 inputs). Our evaluation focuses on the consistency of clothing, identity, and objects. For several practical and well-performed models, we also conducted a user study and evaluated their ELO scores.
As shown in Table 2, our proposed method reports rational scores across several consistency metrics. The Wan+Qwen combo appears to have a leading score in instance score, and this is likely because the image model does not significantly alter or move objects, thus avoiding artifacts that are detected by the VLM question answering. Our method demonstrates comparable scores in object consistency. Furthermore, the user study and ELO scores validate our proposed architecture, confirming that it achieves an effective trade-off between compression and quality.
Different base models. As shown in Table 3, we also evaluate the performance on different base models, focusing on single-frame quality, temporal dynamics, and overall semantic alignment. The results indicate that the Wan-family models achieve high quality on several metrics. Concurrently, larger models, such as the 14B and 12.8B variants, outperform the 5B model in several metrics. The user scores confirm that the 14B Wan model is a well-suited model.
4.5 Alternative Architecture
Figure 8: Adding sliding window. We show that the framework can be combined with a small sliding window to facilitate a continuous shot covering multiple generations.
Table 3: Quantitative results on different base models. We discuss the results using different alternative models as base models. All models are trained as LoRA at rank 128. * The 14B Wan 2.2 is the high noise model.
Frame Temporal Alignment
Method Aesthetic $\uparrow$ Clarity $\uparrow$ Dynamics $\uparrow$ Semantic $\uparrow$ ELO $\uparrow$
HunyuanVideo [36] 12.8B ($\times4$$\times$2) 61.27 67.49 71.22 26.29 1189
Wan [56] 2.2 14B* ($\times4$$\times$2) 67.22 69.37 69.81 27.12 1231
Wan [56] 2.2 5B ($\times4$$\times$2) 66.25 69.01 65.13 25.99 1215
Wan [56] 2.2 5B ($\times2$$\times$2) 66.37 68.95 66.29 26.13 1224
Adding sliding window. As shown in Fig. 8, this method can be combined with a small sliding window (like 3 latent frames) to adjust the occurrence of camera shot shifts. With a sliding window, the frequency of shot changes is reduced in certain cases, allowing the auto-regressive model to generate a continuous shot over multiple generation iterations, resulting in more desirable videos in certain scenarios. For generations without such sliding windows, we recommend generating a complete shot in each generation. Incorporating a sliding window slightly increases the context length.
Figure 9: Adding cross-attention enhancing. We show that connecting the compression model and the DiT with cross-attention layers can improve consistency in difficult cases.
Cross attention enhancement. As shown in Fig. 9, we demonstrate that connecting the features from the penultimate layer of the encoder to each block of the DiT via cross-attention (like IP-Adapter [71]) can further enhance the model’s consistency. This is effective in highly challenging scenarios, e.g., maintaining the arrangement order of items on supermarket shelves. In such detail-oriented situations, adding cross-attention provides a rational consistency boost, at the cost of additional computational costs.
Figure 10: Using multiple memory compression models. We show that using multiple compression models at the same time with different compression patterns, e.g., higher temporal and higher spatial, can facilitate detail consistency in difficult cases, at the cost of doubling the context length.
Combining multiple compression models. As shown in Fig. 10, another enhancement is to use multiple memory encoders simultaneously. For instance, a standard $4\times 4\times 2$ encoder can be paired with a $2\times 2\times 8$ encoder, which prioritizes temporal compression while preserving more spatial details. Using multiple encoders allows the model to capture different aspects of the history. This architecture offers advantages such as preserving fine-grained details, like text on newspapers, magazines, or billboards, at the cost of an increased context length.
5 Conclusion
In this work, we discuss the fundamental trade-off between context length and quality in autoregressive video generation as a critical challenge for creating long, coherent narratives. We introduced a framework pretraining a memory compression model with an explicit objective: high-fidelity retrieval of arbitrary frames from a long video history. This pretraining phase imposes the model to learn a compact yet detail-rich representation of the video context, effectively preserving high-frequency details. By subsequently fine-tuning this pretrained encoder within an autoregressive video diffusion system, we developed a practical and efficient model capable of handling long-term dependencies, such as maintaining character identity and scene consistency.