### v5

# Pretraining Frame Preservation for Lightweight Autoregressive Video History Embedding

Abstract
Autoregressive video generation relies on history context for content consistency and storytelling. As video histories grow longer, efficiently encoding them remains an open problem — particularly for personal users and local workflows where compute and memory budgets are limited. We present a lightweight history encoder that maps long video histories into short-length embeddings, pretrained with a frame query objective that learns to attend to content features at arbitrary temporal positions. The pretraining stage provides the encoder with dense history coverage on large-scale video data; the subsequent finetuning stage adapts the pretrained encoder under an autoregressive video generation objective to establish content-level consistency. In this way, the lightweight embeddings achieve comparable performance to heavier alternatives. We evaluate the framework with ablative settings and discuss the architecture designs.
Figure 1: This video was generated autoregressively second-by-second with full history context (without cutting out any history frames). The full history is encoded into a lightweight embedding and processed by an RTX 4070 12GB.
1 Introduction
Storytelling capability, narrative coherence, and context quality are increasingly important in the video generation community. The latest commercial models like Sora2 [46], Veo3.1 [20], and Kling2.5 [37] enable storyboard creation-based, scene planning, and dynamic camera shots (e.g., in the style of TikTok or YouTube Shorts). Recent academic models have also focused on long video streaming [71, 27], scene planning [22, 26, 82, 85, 86], and video context consistency [28, 31, 42]. Among these approaches, autoregressive models are a key paradigm for video storytelling, supporting native video continuation and storyboard streaming. Autoregressive models view video history as context, and they face unique challenges when handling long-form content. These challenges are amplified for personal users, offline pipelines, and local finetuning, where encoding longer video histories must fit within limited compute and memory.
The trade-off between context quality and length is a typical and critical problem in video context modeling. A naive sliding window, which cuts off all distant-enough frames, maintains a fixed context length but loses long-range history. Various compression approaches have been extensively studied; e.g., highly compressed VAEs like LTXV [23] and DC-AE [10] show that it is possible to compress pixel-space videos into more compact contexts, and hybrid approaches [1, 32, 87] like FramePack [77] show that compression can happen at multiple levels, albeit at the cost of high-frequency image details. Another strategy is to preserve the context length but reduce computation (e.g., using sparse [63, 76, 79, 64] or linear [5, 67, 58, 11, 73, 33] attention), though the linear layer cost remains an overhead for training and bidirectional inference. More specialized methods like token merging [4, 3] also demonstrate that a higher merging rate leads to greater detail loss. These efforts indicate that the balance among history coverage, visual fidelity, and memory cost remains an open problem, and that lightweight history encoding is broadly desired.
Several questions arise in this setting. Should context quality be measured by object, scene, or facial consistency? Do spatial and temporal encoding have different levels of importance? Is there a diminishing return of context length beyond a certain point? A practical approach should address two concerns: learning to represent video history densely and efficiently, and learning to generate content that is consistent with that history. Separating these concerns into a pretraining stage and a finetuning stage reduces the computational cost compared to joint training, and enables each stage to use different data and objectives.
To this end, this paper presents a history encoder that maps video history into a lightweight embedding, learned through a pretraining-then-finetuning approach. In the pretraining stage, the encoder is trained on large-scale video data with a frame query objective: given a long history, the encoder must enable the diffusion model to attend to frames at arbitrary temporal positions. Only a random subset of frames is supervised at each step, ensuring dense history coverage while keeping training cost low. In the finetuning stage, the pretrained encoder is integrated into an autoregressive video diffusion model and adapted jointly to establish content-level consistency.
One challenge in learning such a history encoder is aligning its output with the generation model’s inner feature space. Instead of building the encoded representation from scratch, we reuse the DiT’s own processing pipeline: the encoder outputs directly at the DiT’s inner hidden states (e.g., 3072 or 5120 channels), bypassing the VAE’s narrow bottleneck. This design allows the encoder to manipulate deep DiT features rather than the latent space. Pretraining on millions of videos, this approach provides a strong initialization for the subsequent finetuning stage, reducing training cost and improving the quality of the learned history embedding.
Experiments show that the pretrained encoder can, by itself, query history frames with high feature quality. When integrated into a video diffusion model for autoregressive finetuning, the encoder provides long-range content consistency. This framework leads to a practical autoregressive video model with a lightweight history context and perceptually high consistency, as validated by both quantitative evaluations and human studies.
In summary, we (1) discuss the trade-off between context length and quality in autoregressive video generation, and the need for lightweight history encoding in personal and local workflows; (2) propose to pretrain a history encoder with a frame query objective, providing dense coverage and reducing subsequent finetuning cost; (3) integrate the pretrained encoder into an autoregressive video diffusion system and finetune for content-level consistency; (4) conduct extensive experiments, including quantitative, qualitative, and ablation studies, to analyze the influence of different architecture designs; and (5) provide pretrained models for practical video generation applications.
2 Related Work
Autoregressive video diffusion and long videos. Diffusion has become the driving force behind video synthesis, where a central challenge is length scaling. Training-free length‑extension methods [48, 43, 44, 83, 49] reschedule noise or re-balance temporal frequency to stretch pre-trained models beyond their training horizon. A complementary thread blends diffusion with causal prediction: Diffusion Forcing [7] and HistoryGuidance [52] enable variable-horizon conditioning and stable long rollouts by noise injection. These approaches are adapted in industrial systems such as SkyReels‑V2 [8] and Magi-1 [55]. Additionally, StreamingT2V [24] augments existing models with short- and long-term memory with randomized blending to generate longer videos. FAR [21] studies long‑context AR with flexible RoPE [53] decay and mixed short/long windows, while CausVid [71] distills a bidirectional teacher to a few‑step causal generator. StreamDiT [35] combines multi-step distillation with a moving frame buffer and mixed partition training to generate results in real-time. To mitigate error accumulation during AR generation, Self‑Forcing [27] simulates AR rollout during training, while its extensions [12, 41, 68] further improve length generalization.
Context learning and history encoding. Long-context persistence is crucial as we extend video generation beyond a few seconds. One option is retrieval, which grounds generation in a persistent state. WorldMem [66] and Context-as-Memory [72] augment AR video world models with FoV-based history retrieval, while VMem [39] indexes past views by surfels to retrieve the most relevant viewpoints. Memory Forcing [25] pairs geometry-indexed spatial memory with tailored training regimes to balance exploration and revisits. Beyond fixed retrieval rules, Mixture-of-Contexts [6] learns a dynamic sparse attention context router so that tokens attend only to the most salient chunks. Similarly, MoGA [30] and Holocine [45] also propose token-level sparse attending policies. Pack-and-Force [62] instead proposes a learnable context semantic retriever. Orthogonal to retrieval, compression of history turns an unbounded context into a compact state. FramePack [77] compresses prior frames into a fixed‑size latent “packed” context. Captain Cinema [65] uses a similar compression for keyframes. StateSpaceDiffuser [50] and Po et al. [47] swap quadratic attention for recurrent states to maintain long‑term memory. TTTVideo [13] and LaCT [80] use light MLP test-time training layers as a learned context representation.
Efficient video diffusion designs. As we scale video generation to long horizons, large context windows are bottlenecked, driving a wave of efficient computational designs. Kernel advances such as FlashAttention [14, 15] improve throughput. Static or hardware-friendly sparse patterns include sliding/tiling 3D windows [79], radial spatiotemporal masks [40], and training-free head/pruning heuristics [63, 69]. Dynamic or learned pruning/routing further select salient token pairs or blocks [61], coarse-to-fine sparse token selection [78], Sage/SpargeAttention families [75, 74, 76], blockified routing with cached search [64], and progressive block carving [81]. Another line of work leverages compressing the latent space or token sequence: token merging and patch scaling [2, 38], compact/variable-rate tokenizers [1], highly compressed latent space [23], or multiscale pyramids with re-noising [32]. SANA [67] introduced a linear‑attention Diffusion Transformer for images, while SANA‑Video [9] extends this with block‑linear attention and a constant‑memory KV cache.
Besides, this approach presents distinctions with VAEs: VAEs have bottlenecks with 4, 16, or 64 channels, whereas this method directly works at DiT inner channels like 3072 or 5120 with a different feature scope; VAEs have training overheads to process full contexts, while this method works with lightweight partial overheads (frame query objective).
3 Method
We pretrain a history encoder with the objective to query content features at arbitrary time positions (Section 3.1). The pretrained encoder is then finetuned within autoregressive video diffusion models (Section 3.2).
Preliminaries. We clarify the notation of a typical autoregressive video diffusion model. Unless otherwise noted, all “frames”, “pixels”, etc., refer to latent concepts. We consider the typical Diffusion Transformers (DiTs) like Wan [56] and HunyuanVideo [36] with flow match scheduling. Rectified-flow models map noisy latents $\bm{X}_{t_{i}}\in\mathbb{R}^{T\times H\times W\times C}$ with
$\bm{X}_{t_{i}}=(1-t_{i})\bm{X}_{0}+t_{i}\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}\big(\mathbf{0},\mathbf{I}\big),$ (1)
from clean latents $\bm{X}_{0}$ where $t_{i}\in(0,1]$ is the diffusion timestep. The autoregressive model typically takes the history $\bm{H}\in\mathbb{R}^{T_{h}\times H\times W\times C}$ as conditioning. The generator $\bm{G}_{\theta}(\cdot)$ is trained by finding
$\begin{split}\arg\min_{\theta}&\mathbb{E}_{\bm{X}_{0},\bm{H},\bm{c},\bm{\epsilon},t_{i}\sim\mathcal{L}(0,1)}\\ &\bigg\|(\bm{\epsilon}-\bm{X}_{0})-\bm{G}_{\theta}\big(\bm{X}_{t_{i}},t_{i},\bm{c},\bm{H}\big)\bigg\|_{2}^{2},\end{split}$ (2)
where $\bm{c}$ is a set of conditions such as text prompts (a single prompt or streamed from a storyboard), and $t_{i}\sim\mathcal{L}(0,1)$ is the shifted logit-normal distribution [18] for flow matching.
3.1 Pretraining History Encoder
Figure 2: Pretraining the history encoder. The encoder maps long video histories into lightweight embeddings. The objective of the pretraining is to query content features at arbitrary history time positions.
Consider a typical case where we have a history of a 60-second video and want to generate the next 1 second, in a typical setting of 480p, 24fps using standard Hunyuan/Wan VAE and patchifying. The history context length becomes $832/16\times 480/16\times 60\times 24/4=561,600$, which creates an inference bottleneck for consumer-level devices that have limited GPU memory (and a training bottleneck for clusters even with causal computation or KV-cache designs). A naive sliding window will fail to model long-range consistency. Typical overhead reduction like token merging, downsampling, larger patch sizes, compact VAEs, etc., come with costs of losing image detail at various levels.
We observe that a useful indicator of the history encoder’s ability to preserve content features is its capacity to query frames at arbitrary temporal positions. Given a video history $\bm{H}$, we randomly sample a set of frame indices $\bm{\Omega}$. Frames at the selected indices are kept clean, while all remaining frames are masked with noise sampled from $\mathcal{L}(0.2,1)$, resulting in a masked history $\texttt{mask}(\bm{H},\bm{\Omega})$. As shown in Fig. 2, the DiT learns to query the selected frames via a conditional mapping
$\bm{G}_{\theta}:\{\phi(\texttt{mask}(\bm{H},\bm{\Omega})),\bm{c}\}\mapsto\bm{H}_{[\bm{\Omega}]},$ (3)
trained by finding
$\begin{split}&\arg\min_{\theta,\phi}\mathbb{E}_{\bm{H},\bm{\Omega},\bm{c},\bm{\epsilon},t_{i}}\\ &\bigg\|(\bm{\epsilon}-\bm{H}_{[\bm{\Omega}]})-\bm{G}_{\theta}\big((\bm{H}_{[\bm{\Omega}]})_{t_{i}},t_{i},\bm{c},\phi(\texttt{mask}(\bm{H},\bm{\Omega}))\big)\bigg\|_{2}^{2},\end{split}$ (4)
where $\bm{G}_{\theta}(\cdot)$ is the DiT with low-rank adaptation (LoRA).
When the output of the encoder $\phi(\cdot)$ is short relative to the original history, the frame query task becomes highly difficult, making it well-suited for pretraining history encoders on millions of in-the-wild videos. The randomness of $\bm{\Omega}$ is important: if only frames at fixed positions (e.g., the last few frames) are used as targets, the encoder can learn a cheating shortcut by dedicating all capacity to those positions while ignoring the rest of the history. Random selection forces the embedding to maintain dense coverage across the entire temporal extent. Besides, directly training to query all frames is impractical with minute-level data, even for highly efficient infrastructures.
Figure 3: Architecture of the history encoder. We use 3D convolution, SiLU, and attention to establish a lightweight neural structure as the baseline encoder. Different alternative architectures (e.g., various channels, full transformer, etc.) are possible and will be discussed in ablation.
Network architecture. As shown in Fig. 3, we present a lightweight baseline architecture for the history encoder $\phi(\cdot)$. Instead of building the compressed representation from scratch, we reuse the DiT’s context representation: we first downsample a high-resolution high-fps video into a low-resolution low-fps one and process it with the VAEs and DiT’s patchifiers and first-layer projections. After that, we encode the high-resolution original video into a residual-enhancing vector and add it to the context vector. Note that the feature addition happens after the DiT’s first-layer projection, which means the encoder does not go through the VAE’s 16 channel bottleneck but directly outputs at inner channels (like 3072 for WAN-5B) for best fidelity. The baseline encoder uses 3D convolutions as starting layers and ends with attention layers. Besides, it is also possible to connect the DiT with cross-attentions from the last hidden states of the encoder for another refinement pass, but the quality gain is marginal and will be covered in ablations.
3.2 Finetuning Video Diffusion Model
Figure 4: Finetuning autoregressive video models. We illustrate the finetuning and inference of the final autoregressive video models. The pretraining of the history encoder is finished before the finetuning.
With the pretrained history encoder $\phi(\cdot)$, we finetune the video diffusion model (with LoRA) together with the encoder for autoregressive generation. Unlike the pretraining stage, the history $\bm{H}$ is now complete and noise-free, and no frame indices $\bm{\Omega}$ are involved. The DiT learns to produce the next video segment $\bm{X}$ conditioned on the history embedding
$\bm{G}_{\theta}:\{\phi(\bm{H}),\bm{c}\}\mapsto\bm{X},$ (5)
by the diffusion objective
$\begin{split}\arg\min_{\theta,\phi}&\mathbb{E}_{\bm{X}_{0},\bm{H},\bm{c},\bm{\epsilon},t_{i}}\\ &\bigg\|(\bm{\epsilon}-\bm{X}_{0})-\bm{G}_{\theta}\big(\bm{X}_{t_{i}},t_{i},\bm{c},\phi(\bm{H})\big)\bigg\|_{2}^{2},\end{split}$ (6)
where newly initialized LoRA parameters are applied. As shown in Fig. 4-(a), the encoder $\phi(\cdot)$ continues to be updated during finetuning, allowing its features to adapt from frame-level query targets toward generation-oriented consistency. The inference follows autoregressive generation by repeatedly concatenating the generated sections into the history as shown in Fig. 4-(b). Since the encoder is almost fully convolutional, the encoded history can also be concatenated on-the-fly without recomputation.
Figure 5: Visual comparison on frame query results. We present results after the pretraining using different possible neural structures and various compression settings. * The “Large Patchifier” is technically equivalent to [77].
4 Experiment
4.1 Experimental Details
[…]
Base models. All experiments are conducted on base models of HunyuanVideo [36] and Wan [56] family. For Wan 2.2, the involved models are the 5B model and the 14B high noise model. Only the high noise model is used to simplify experiments (similar to the design choice of WanAnimate).
Data preparation. The dataset is prepared with about 5 million internet videos from a range of websites. About half of the data is vertical Short-style videos and the rest are common horizontal videos. The data is cleaned for quality, and then the high-quality part is captioned with Gemini-2.5-flash VLM and the remainder is processed with local VLMs like QwenVL [57]. The generated captions are in storyboard format with timestamps. When training the autoregressive models, only the prompt with the nearest timestamp is received by the model.
Storyboards. If not noticed, all storyboard prompts in this paper are written by Gemini-2.5-pro. We also tested open source alternatives like Qwen [57] and the quality difference is negligible. See also the supplement for instructions.
Test samples. The test set includes 1000 storyboard prompts written by Gemini-2.5-pro and 4096 unseen videos independent from the training dataset.
4.2 Ablation Study
Table 1: Quantitative results on encoder architectures. We show numerical tests with different ablative architectures. * The “Large Patchifier” is technically equivalent to [77].
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Large Pachifier* ($\times4$$\times$2) 12.93 0.412 0.365
Only LR ($\times4$$\times$2) 15.21 0.472 0.212
Without LR ($\times4$$\times$2) 15.73 0.423 0.198
Proposed ($\times4$$\times$2) 17.41 0.596 0.171
Proposed ($\times2$$\times$2) 19.12 0.683 0.152
Proposed ($\times2$$\times$4) 18.63 0.637 0.153
Proposed ($\times2$$\times$1) 20.19 0.705 0.121
Encoder architecture. As shown in Fig. 5, we compare the image reconstruction details of different architectures under various compression rates, focusing on the performance of Only LR, Without LR, Large Patchifier, and our proposed method. The Only LR variant keeps only the low-resolution branch in Fig. 3, while Without LR removes it, leaving only the high-resolution branch. The Large Patchifier method increases the compression rate by enlarging the DiT’s patchifying projection kernel without introducing new layers, equivalent to FramePack [77] and we directly use its implementation. The results indicate that the Only LR and Without LR variants produce reconstructions that differ significantly from the original images. The Large Patchifier method yields substantial structural changes. The proposed method, even at a high compression rate such as $4\times 4\times 2$, can still preserve much of the image appearance. A lower compression rate, such as $2\times 2\times 1$, provides the best detail preservation, albeit at the cost of a longer context length.
The quantitative results are presented in Table 1. These results show that this method achieves relatively higher performance on metrics like PSNR and SSIM. Furthermore, even at a higher compression rate of $4\times 4\times 2$, this method effectively preserves the original image structure.
Figure 6: Influence of history encoder pretraining. We present results with or without the pretraining of the history encoder. The input is the same history video, and we visualize the middle frames in the output shots.
Influence of pretraining. We demonstrate in Fig. 6 the impact of pre-training by comparing models trained with and without it. We keep the model architecture and the 20-second history input identical. One model is trained from randomly initialized encoder weights, while the other is fine-tuned using the pretrained memory encoder. Both models are trained with 100k sufficient steps. The results show a marked difference. With pre-training, the model maintains strong temporal consistency, leading to notable improvements in facial features, clothing, global video style, storytelling plots, and the coordination of camera movements. In contrast, the model trained without pre-training sometimes fails to attend to relevant frames in the history, resulting in inconsistencies.
Error accumulation (drifting). The problem of error accumulation highly depends on the training dataset. Surprisingly, we find that when trained on videos with dense camera shot shifts, like Short-style videos, the error accumulation problem seems to be absent or invisible. For other types of videos requiring long continuation over single shots, multiple design choices [77, 27, 7] are discussed in the supplement.
4.3 Qualitative Results
Figure 7: Qualitative results on storyboards. We present results by streaming prompts from storyboards. A storyboard is a set of prompts where each prompt covers a certain number of frames. The storyboards can be written by external language models.
Storyboard results. As shown in Fig. 7, we demonstrate that our model can handle a diverse range of prompts and storyboards while maintaining consistency across characters, scenes, objects, and plotlines.
4.4 Quantitative Results
Metrics. We include several metrics for video evaluations from VBench [29], VBench2 [84], etc., with some modifications. Cloth: The cloth consistency by asking VLM questions. We upgrade VBench2 [84]’s LLaVA to Gemini-2.5-pro; Instance: The object-level quality by asking VLM questions about unnatural or inconsistent instances. We upgrade VBench2 [84]’s Qwen to Gemini-2.5-pro; Identity: The facial similarity using ArcFace [16] with face detection RetinaFace [17]; Clarity: The MUSIQ [34] image quality predictor trained on SPAQ [19]. This metric evaluates artifacts like noises; Aesthetic: The LAION aesthetic predictor [51], measuring the aesthetic values perceived by a CLIP estimator; Dynamic: The RAFT [54] modified by VBench to estimate the degree of dynamics. Note that multi-shot videos intuitively have a special RAFT flow between shots, but in practice, RAFT seems to output near-zero values between shots, making the difference negligible from regular videos; Semantic: The video-text score computed by ViCLIP [59]. This metric measures the overall semantic consistency and prompt alignments.
Table 2: Quantitative results on video content consistency. We present the performance of different possible architectures on content consistency. The “1p” in Qwen means using 1 image as an image model input. Methods with severe artifacts are excluded from human ELO scores.
Human Object User Study
Method Cloth $\uparrow$ Identity $\uparrow$ Instance $\uparrow$ ELO $\uparrow$
WanI2V [56] + QwenEdit [60] (1p) 94.10 68.45 85.21 /
WanI2V [56] + QwenEdit [60] (2p) 95.09 68.22 91.19 1198
WanI2V [56] + QwenEdit [60] (3p) 94.28 67.30 80.34 /
Only LR ($\times4$$\times$2) 91.98 69.22 85.32 1194
Without LR ($\times4$$\times$2) 89.64 67.41 82.85 /
Without Pretrain ($\times4$$\times$2) 87.12 66.99 81.13 /
Proposed ($\times4$$\times$2) 96.12 70.73 89.89 1216
Proposed ($\times2$$\times$2) 96.71 72.12 90.27 1218
Consistency analysis. We compare various architectures of our method, including different compression rates and ablations with or without pre-training and the LR branch. We also compare against a common baseline that combines Wan-2.2-I2V [56] with Qwen-Image-Edit-2509 [60]. This baseline first generates a series of initial frames with an image model, animates them into video clips, and then concatenates the clips. We test this baseline with 1, 2, and 3 images for Qwen-Edit as history contexts (since the model supports at most 3 inputs). Our evaluation focuses on the consistency of clothing, identity, and objects. For several practical and well-performed models, we also conducted a user study and evaluated their ELO scores.
As shown in Table 2, our proposed method reports rational scores across several consistency metrics. The Wan+Qwen combo appears to have a leading score in instance score, and this is likely because the image model does not significantly alter or move objects, thus avoiding artifacts that are detected by the VLM question answering. Our method demonstrates comparable scores in object consistency. Furthermore, the user study and ELO scores validate our proposed architecture, confirming that it achieves an effective trade-off between compression and quality.
Different base models. As shown in Table 3, we also evaluate the performance on different base models, focusing on single-frame quality, temporal dynamics, and overall semantic alignment. The results indicate that the Wan-family models achieve high quality on several metrics. Concurrently, larger models, such as the 14B and 12.8B variants, outperform the 5B model in several metrics. The user scores confirm that the 14B Wan model is a well-suited model.
4.5 Alternative Architecture
Figure 8: Adding sliding window. We show that the framework can be combined with a small sliding window to facilitate a continuous shot covering multiple generations.
Table 3: Quantitative results on different base models. We discuss the results using different alternative models as base models. All models are trained as LoRA at rank 128. * The 14B Wan 2.2 is the high noise model.
Frame Temporal Alignment
Method Aesthetic $\uparrow$ Clarity $\uparrow$ Dynamics $\uparrow$ Semantic $\uparrow$ ELO $\uparrow$
HunyuanVideo [36] 12.8B ($\times4$$\times$2) 61.27 67.49 71.22 26.29 1189
Wan [56] 2.2 14B* ($\times4$$\times$2) 67.22 69.37 69.81 27.12 1231
Wan [56] 2.2 5B ($\times4$$\times$2) 66.25 69.01 65.13 25.99 1215
Wan [56] 2.2 5B ($\times2$$\times$2) 66.37 68.95 66.29 26.13 1224
Adding sliding window. As shown in Fig. 8, this method can be combined with a small sliding window (like 3 latent frames) to adjust the occurrence of camera shot shifts. With a sliding window, the frequency of shot changes is reduced in certain cases, allowing the auto-regressive model to generate a continuous shot over multiple generation iterations, resulting in more desirable videos in certain scenarios. For generations without such sliding windows, we recommend generating a complete shot in each generation. Incorporating a sliding window slightly increases the context length.
Figure 9: Adding cross-attention enhancing. We show that connecting the history encoder and the DiT with cross-attention layers can improve consistency in difficult cases.
Cross attention enhancement. As shown in Fig. 9, we demonstrate that connecting the features from the penultimate layer of the encoder to each block of the DiT via cross-attention (like IP-Adapter [70]) can further enhance the model’s consistency. This is effective in highly challenging scenarios, e.g., maintaining the arrangement order of items on supermarket shelves. In such detail-oriented situations, adding cross-attention provides a rational consistency boost, at the cost of additional computational costs.
Figure 10: Using multiple history encoders. We show that using multiple encoders at the same time with different compression patterns, e.g., higher temporal and higher spatial, can facilitate detail consistency in difficult cases, at the cost of doubling the context length.
Combining multiple encoders. As shown in Fig. 10, another enhancement is to use multiple history encoders simultaneously. For instance, a standard $4\times 4\times 2$ encoder can be paired with a $2\times 2\times 8$ encoder, which prioritizes temporal compression while preserving more spatial details. Using multiple encoders allows the model to capture different aspects of the history. This architecture offers advantages such as preserving fine-grained details, like text on newspapers, magazines, or billboards, at the cost of an increased context length.
5 Conclusion
In this work, we study the challenge of encoding long video histories into lightweight context for autoregressive video generation, targeting personal and local workflows with limited compute. We present a history encoder pretrained with a frame query objective: the encoder learns to attend to arbitrary frames from a long video history and encode them into a lightweight embedding. By subsequently finetuning this pretrained encoder within an autoregressive video diffusion system, the resulting model maintains long-range content consistency, including character identity, clothing, and scene layout. We evaluate the framework with ablative settings and discuss the trade-offs of different architecture designs.